# Superconductivity in twisted WSe2 from topology-induced quantum fluctuations

### v1

@@ Acknowledgements.
Table 1: Parameters of the single valley tight-binding model with $\theta=1.43^{\circ}$ and $\varepsilon_{D}=3\,\rm meV$. The variables $R_{1},R_{2}\in\mathbb{Z}$ stand for the components of the hopping vector $\mathbf{R}=R_{1}\mathbf{a}_{1}+R_{2}\mathbf{a}_{2}$. Note that the first two rows are the on-site energies of the two orbitals.
Appendix A Continuum model
In this appendix, we provide a brief discussion about the continuum model of twisted homo-bilayer TMDC, which was introduced in Ref. [24]. With the inter-layer and intra-layer moiré potential considered, the low-energy effective Hamiltonian of twisted bilayer WSe2 can be written as:
$H=\left(\begin{array}[]{cc}\frac{\nabla^{2}}{2m^{*}}+v_{+}(\mathbf{r})+\frac{\varepsilon_{D}}{2}&t(\mathbf{r})\\ t^{*}(\mathbf{r})&\frac{\nabla^{2}}{2m^{*}}+v_{-}(\mathbf{r})-\frac{\varepsilon_{D}}{2}\end{array}\right)\,.$ (7)
in which $m^{*}$ is the effective mass of the hole pocket, $v_{\pm}(\mathbf{r})$ is the moiré potential in the top and bottom layers, respectively, and $t(\mathbf{r})$ is the inter-layer tunneling term. The displacement field potential $\varepsilon_{D}$ is also considered in the Hamiltonian. In the absence of the displacement field, the Hamiltonian Eq. (7) possess both $C_{3z}$ and $C_{2y}T$ symmetries, while the displacement field term will break the $C_{2y}T$ symmetry. Under the lowest-harmonic approximation [24], the $C_{3z}$ symmetry enforces the intra-layer periodic potential to have the following form:
$v_{\ell}(\mathbf{r})=2\tilde{v}\sum_{j=1,3,5}\cos(\mathbf{g}_{j}\cdot\mathbf{r}+\ell\psi)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \ell=\pm\,,$ (8)
$t(\mathbf{r})=w\sum_{j=1,2,3}e^{i{\mathbf{q}}_{j}\cdot\mathbf{r}}\,,$ (9)
in which the definition of the three ${\mathbf{q}}_{j}$ vectors can be found in Fig. 1(b). By fitting the continuum Hamiltonian Eq. (7) with ab initio simulations, the parameters for twisted bilayer $\rm WSe_{2}$ are determined to be $a_{0}=3.317\,\text{\AA}$, $m^{*}=0.43m_{e}$, $\tilde{v}=9\,{\rm meV}$, $\psi=128^{\circ}$ and $w=18\,\rm meV$, in which $m_{e}$ is the mass of the free electron [25]. Band structure shown in Fig. 2 is computed from this continuum model.
Appendix B Band topology
In this appendix, we will discuss the band structure and the associated band topology of twisted bilayer WSe2 with different moiré potential strengths. As indicated by the continuum model Hamiltonian in Eq. (7), the band structure is mostly determined by the twisting angle $\theta$ and the moiré potential strength $\tilde{v}$ and $w$, and is expected to be sensitive to such parameters [54].
Here, we study the band structure and the topology of the top bands in the case of tWSe2 with the angle $\psi=128^{\circ}$ and effective mass $m^{*}=0.43m_{e}$ fixed. In Fig. 5, we show the direct band gaps between the first and the second top-most bands $\Delta_{12}$, and between the second and the third top-most bands $\Delta_{23}$ as functions of the moiré potential strength $\tilde{v}$ and $w$, with the twisting angles $\theta=3.65^{\circ}$ and $\theta=1.43^{\circ}$. We vary the strength of the intra-layer potential $\tilde{v}$ from $5$ to $15\,\rm meV$, and the inter-layer potential $w$ from $10$ to $30\,\rm meV$. It can be seen that with a moderate change of the moiré potential strength, it is possible to obtain a band structure with the top two bands carrying zero total Chern number, even if the twisting angle is $\theta=3.65^{\circ}$.
Appendix C Tight-binding parameters
The values of the hopping parameters for the two-orbital tight-binding model, which are greater than $0.1\,\rm meV$, with parameters $\theta=1.43^{\circ}$ and $\varepsilon_{D}=3\,\rm meV$, are listed in Table 1.

### v2

@@ C Tight-binding parameters
Table S1: Parameters of the single valley tight-binding model with $\theta=1.43^{\circ}$ and $\varepsilon_{D}=3\,\rm meV$. The variables $R_{1},R_{2}\in\mathbb{Z}$ stand for the components of the hopping vector $\mathbf{R}=R_{1}\mathbf{a}_{1}+R_{2}\mathbf{a}_{2}$. Note that the first two rows are the on-site energies of the two orbitals.
$1$ $2$ $2$ 00 00 13.76 $180.0^{\circ}$
$2$ $2$ $2$ $-1$ $-1$ 4.38 $-115.3^{\circ}$
$3$ $2$ $2$ $-1$ 00 4.38 $115.3^{\circ}$
$4$ $2$ $2$ 00 $-1$ 4.38 $115.3^{\circ}$
$5$ $2$ $2$ 00 $1$ 4.38 $-115.3^{\circ}$
$6$ $2$ $2$ $1$ 00 4.38 $-115.3^{\circ}$
$7$ $2$ $2$ $1$ $1$ 4.38 $115.3^{\circ}$
$8$ $1$ $1$ $-1$ $-1$ 4.25 $118.2^{\circ}$
$9$ $1$ $1$ $-1$ 00 4.25 $-118.2^{\circ}$
$10$ $1$ $1$ 00 $-1$ 4.25 $-118.2^{\circ}$
$11$ $1$ $1$ 00 $1$ 4.25 $118.2^{\circ}$
$12$ $1$ $1$ $1$ 00 4.25 $118.2^{\circ}$
$13$ $1$ $1$ $1$ $1$ 4.25 $-118.2^{\circ}$
$14$ $1$ $2$ $-1$ 00 3.51 $-48.6^{\circ}$
$15$ $2$ $1$ 00 $-1$ 3.51 $48.6^{\circ}$
$16$ $2$ $1$ 00 00 3.51 $48.6^{\circ}$
$17$ $1$ $2$ 00 00 3.51 $-48.6^{\circ}$
$18$ $1$ $2$ 00 $1$ 3.51 $-48.6^{\circ}$
$19$ $2$ $1$ $1$ 00 3.51 $48.6^{\circ}$
$20$ $2$ $1$ $-1$ $-1$ 1.76 $-131.4^{\circ}$
$21$ $1$ $2$ $-1$ $-1$ 1.76 $131.4^{\circ}$
$22$ $1$ $2$ $-1$ $1$ 1.76 $131.4^{\circ}$
$23$ $2$ $1$ $1$ $-1$ 1.76 $-131.4^{\circ}$
$24$ $2$ $1$ $1$ $1$ 1.76 $-131.4^{\circ}$
$25$ $1$ $2$ $1$ $1$ 1.76 $131.4^{\circ}$
$26$ $1$ $1$ 00 00 0.93 $180.0^{\circ}$
Table S2: Parameters of the single valley tight-binding model with $\theta=3.65^{\circ}$ and $\varepsilon_{D}=15\,\rm meV$. The moiré potential strength in this calculation is chosen to be $\tilde{v}=13\rm\,meV$ and $w=15\,\rm meV$.
D Projected Coulomb interactions
Except for the tight-binding parameters shown in the previous section, another output of Wannier90 software is the unitary gauge transformation $U_{\alpha,n}({\mathbf{k}})$, which rotates the single valley eigenstates with the form of Eq. (S4), into the Bloch states that correspond to the two local orbitals:
$\psi_{\alpha{\mathbf{k}}}(\mathbf{r},\ell)$ $=\frac{1}{\sqrt{\Omega_{\rm tot}}}\sum_{n={1,2}}U_{\alpha,n}({\mathbf{k}})\sum_{\mathbf{Q}\in\mathcal{Q}_{\ell}}u_{\mathbf{Q},\alpha}({\mathbf{k}})e^{i({\mathbf{k}}-\mathbf{Q})\cdot\mathbf{r}}=\frac{1}{\sqrt{\Omega_{\rm tot}}}\sum_{\mathbf{Q}\in\mathcal{Q}_{\ell}}\tilde{u}_{\mathbf{Q},\alpha}({\mathbf{k}})e^{i({\mathbf{k}}-\mathbf{Q})\cdot\mathbf{r}}\,,$ (S8)
$\tilde{u}_{\mathbf{Q},\alpha}({\mathbf{k}})$ $=\sum_{n=1,2}U_{\alpha,n}({\mathbf{k}})u_{\mathbf{Q},n}({\mathbf{k}})\,.$ (S9)
In general, the interacting Hamiltonian can be written as the following form:
$H_{I}$ $=\frac{1}{2\Omega_{\rm tot}}\sum_{{\mathbf{k}}{\mathbf{k}}^{\prime}{\mathbf{q}}}\sum_{\alpha\beta\alpha^{\prime}\beta^{\prime}\sigma\sigma^{\prime}}U_{\alpha\beta\alpha^{\prime}\beta^{\prime}}^{(\sigma\sigma^{\prime})}({\mathbf{q}};{\mathbf{k}},{\mathbf{k}}^{\prime})c^{\dagger}_{{\mathbf{k}}+{\mathbf{q}}\alpha\sigma}c_{{\mathbf{k}}\beta\sigma}c^{\dagger}_{{\mathbf{k}}^{\prime}-{\mathbf{q}}\alpha^{\prime}\sigma^{\prime}}c_{{\mathbf{k}}^{\prime}\beta^{\prime}\sigma^{\prime}}\,,$ (S10)
$U^{(\sigma\sigma^{\prime})}_{\alpha\beta\alpha^{\prime}\beta^{\prime}}({\mathbf{q}};{\mathbf{k}},{\mathbf{k}}^{\prime})$ $=\sum_{\mathbf{G}\in\mathcal{Q}_{0}}V({\mathbf{q}}+\mathbf{G})\sum_{\mathbf{Q}\in\mathcal{Q}_{\pm}}\tilde{u}^{(\sigma)*}_{\mathbf{Q},\alpha}({\mathbf{k}}+{\mathbf{q}}+\mathbf{G})\tilde{u}^{(\sigma)}_{\mathbf{Q},\beta}({\mathbf{k}})\sum_{\mathbf{Q}\in\mathcal{Q}_{\pm}}\tilde{u}^{(\sigma^{\prime})*}_{\mathbf{Q}^{\prime},\alpha^{\prime}}({\mathbf{k}}^{\prime}-{\mathbf{q}}-\mathbf{G})\tilde{u}^{(\sigma^{\prime})}_{\mathbf{Q}^{\prime},\beta^{\prime}}({\mathbf{k}}^{\prime})\,.$ (S11)
Here $\tilde{u}^{(\uparrow)}_{\mathbf{Q},\alpha}({\mathbf{k}})$ are the Bloch state orbital wave functions shown in Eq. (S9), and $\tilde{u}^{(\downarrow)}_{\mathbf{Q},\alpha}({\mathbf{k}})$ are generate from $\tilde{u}^{(\uparrow)}_{\mathbf{Q},\alpha}({\mathbf{k}})$ through time-reversal transformation. The Fourier transformation of the screened Coulomb interaction is given by the form of $V({\mathbf{q}})=(\xi e^{2}/4\varepsilon_{0}\varepsilon)\tanh(\xi q/2)/(\xi q/2)$, where $\xi$ is the gate distance, and $\varepsilon$ is the relative dielectric constant. This interaction vertex $U^{(\sigma\sigma^{\prime})}_{\alpha\beta\alpha^{\prime}\beta^{\prime}}({\mathbf{q}};{\mathbf{k}},{\mathbf{k}}^{\prime})$ can be Fourier transformed into the real space orbitals as:
$H_{I}=$ $\frac{1}{2}\sum_{\mathbf{R}\mathbf{d}\mathbf{d}^{\prime}\mathbf{R}_{0}}\sum_{\alpha\beta\alpha^{\prime}\beta^{\prime}\sigma\sigma^{\prime}}\tilde{U}^{(\sigma\sigma^{\prime})}_{\alpha\beta\alpha^{\prime}\beta^{\prime}}(\mathbf{R};\mathbf{d},\mathbf{d}^{\prime})c^{\dagger}_{\mathbf{R}+\mathbf{d}+\mathbf{R}_{0}\alpha\sigma}c_{\mathbf{d}+\mathbf{R}_{0}\beta\sigma}c^{\dagger}_{\mathbf{d}^{\prime}+\mathbf{R}_{0}\alpha^{\prime}\sigma^{\prime}}c_{\mathbf{R}_{0}\beta^{\prime}\sigma^{\prime}}\,,$ (S12)
$\tilde{U}^{(\sigma\sigma^{\prime})}_{\alpha\beta\alpha^{\prime}\beta^{\prime}}(\mathbf{R;\mathbf{d},\mathbf{d}^{\prime}})=$ $\frac{1}{N^{2}\Omega_{\rm tot}}\sum_{{\mathbf{k}}{\mathbf{k}}^{\prime}{\mathbf{q}}}U_{\alpha\beta\alpha^{\prime}\beta^{\prime}}^{(\sigma\sigma^{\prime})}({\mathbf{q}};{\mathbf{k}},{\mathbf{k}}^{\prime})e^{i{\mathbf{q}}\cdot(\mathbf{R}+\mathbf{d}-\mathbf{d}^{\prime})}e^{i{\mathbf{k}}\cdot\mathbf{d}}e^{i{\mathbf{k}}^{\prime}\cdot\mathbf{d}^{\prime}}\,.$ (S13)
We computed the value of the projected screened Coulomb interaction with gate distance $\xi=3\,\rm nm$ at $\mathbf{R}=\mathbf{d}=\mathbf{d}^{\prime}=\mathbf{0}$ (on-site interaction) for both orbitals with different model parameters. Using the parameters $(\tilde{v},w,\varepsilon_{D},\theta)=(9{\rm\,meV},18{\rm\,meV},3{\rm\,meV},1.43^{\circ})$, we obtain the interaction strengths as $\varepsilon U_{1}=182.25\,\rm meV$ and $\varepsilon U_{2}=154.35\rm\,meV$. Using the parameters $(\tilde{v},w,\varepsilon_{D},\theta)=(13{\rm\,meV},15{\rm\,meV},15{\rm\,meV},3.65^{\circ})$, we obtain the interaction strength $\varepsilon U_{1}=520.5\rm\,meV$ and $\varepsilon U_{2}=468.15\rm\,meV$.
Since the tWSe2 sample is sandwiched between h-boron-nitride substrates, the reasonable estimation for the effective dielectric function should be in between $\varepsilon_{\rm hBN}\approx 6$ and $\varepsilon_{\rm WSe_{2}}\approx 16$ [65], which will provide the estimation values for on-site interactions upper and lower limits, respectively. In the case with twisting angle $\theta=1.43^{\circ}$, the interaction strength will be $10\sim 30\,\rm meV$, while in the case with twisting angle $\theta=3.65^{\circ}$, the interaction strength is around $30\sim 85\,\rm meV$.
E Extended Dynamical mean-field theory
In this section, we describe the numerical method used to solve the many-body Hamiltonian. As discussed in the main text, when a finite potential difference is applied, the filling factors of the two orbitals in the effective model differ significantly. One orbital is nearly half-filled, while the other is not. Consequently, despite similar on-site interaction strengths, the two orbitals exhibit distinct localization tendencies. This difference creates a parameter regime analogous to that of a heavy-fermion model.
To simplify the calculation, we explicitly include the emergent RKKY exchange interactions between the local moments. By tuning the strength of the RKKY interaction, a Kondo destruction quantum critical point can be realized. This QCP arisesin the framework of EDMFT, in which the single-particle Green’s function, the spin susceptibility and the local spin moment expectation values are solved in a self-consistent manner [38, 39]. More precisely, we embed the $f$ orbital into a fermion bath $G^{-1}_{0}(i\omega)$ and a bosonic bath $\chi^{-1}_{0}(i\omega)$ to mimic the effect of the hybridization with conduction electrons and the RKKY exchange. The $f$ orbital Green’s function, self energy $\Sigma(i\omega)$ and the fermionic bath $G^{-1}_{0}(i\omega)$ are determined via the following self-consistent condition:
$G_{\rm loc}(i\omega)$ $=\frac{1}{N}\sum_{{\mathbf{p}}}\frac{1}{i\omega-\epsilon_{f}-\Sigma(i\omega)-\frac{V_{\rm hyb}^{2}}{i\omega-\epsilon_{\mathbf{p}}}}\,,$ (S14)
$\Sigma(i\omega)$ $=G^{-1}_{0}(i\omega)-G_{\rm cl}^{-1}(i\omega)\,,$ (S15)
$G_{\rm cl}(i\omega)$ $=G_{\rm loc}(i\omega)\,.$ (S16)
Similarly, the spin susceptibility and the bosonic bath $\chi_{0}^{-1}(i\omega)$ are also determined by another set of self-consistent conditions:
$\chi_{\rm cl}^{-1}(i\omega)$ $=\chi_{0}^{-1}(i\omega)+M(i\omega)\,,$ (S17)
$\chi_{\rm loc}(i\omega)$ $=\frac{1}{N}\sum_{{\mathbf{q}}}\frac{1}{I_{{\mathbf{q}}}+M(i\omega)}\,,$ (S18)
$\chi_{\rm loc}(i\omega)$ $=\chi_{\rm cl}(i\omega)\,,$ (S19)
in which the cluster Green’s function $G_{\rm cl}(i\omega)$ and the susceptibility $\chi_{\rm cl}(i\omega)$ are solved exactly accounting for the effects of the fermionic and bosonic baths. $I_{\mathbf{q}}$ is the Fourier transformation of the RKKY interaction, which controls the controls the competition between the Kondo screening and the tendency towards ordering states of the local moments. In addition to the cluster Green’s function and susceptibility, the local spin moment expectation value can also be evaluated from the quantum Monte Carlo simulation. Therefore, by solving the self-consistent EDMFT equations with different RKKY interaction strengths and at different temperatures, one can determine the critical value of the exchange strength $I_{c}$, and ordering transition temperature $T_{\rm order}$. We also note that the quantum Monte Carlo simulation is conducted using imaginary frequencies, requiring an additional analytical continuation step to obtain the spectral function, which is in general a difficult procedure. Methods that work directly in the real frequency are necessary to properly address this issue, and we leave it for future study.
To allow for the development of unconventional pairing, we adopt a cluster version of EDMFT (C-EDMFT), with the cluster containing two sites of $f$-orbitals that are coupled to each other by the antiferromagnetic RKKY interaction. The details of the approach are given in the literature [40, 47]. In the following, we focus on the spin-singlet pairing that is driven by the antiferromagnetic RKKY interaction. The static pairing susceptibility (at zero wavevector), $\chi_{SC}(T)$, involves a form factor, which, for our case of a triangular lattice, is given by $f({\mathbf{k}})=\cos k_{1}+\cos k_{2}+\cos k_{3}$, where $k_{i}={\mathbf{k}}_{i}\cdot\mathbf{a}_{i}$, with $\mathbf{a}_{3}=-\mathbf{a}_{1}-\mathbf{a}_{2}$. The pairing susceptibility is determined by a Bethe Salpeter equation, in terms of a bare particle-particle bubble and a irreducible paring vertex function [40]. The Brillouin zone is effectively divided into two distinct regions. For a triangular lattice, we select the cluster momenta as ${\mathbf{K}}_{+}=(0,0)$ and ${\mathbf{K}}_{-}=(\pm)\frac{1}{3}(\mathbf{b}_{1}+\mathbf{b}_{2})$, and ${\mathbf{K}}$(${\mathbf{K}}^{\prime}$) represents either ${\mathbf{K}}_{\pm}$. The superconducting phase is identified by the divergence of $\chi_{SC}(T)$, which specifies the superconducting transition temperature $T_{c}$.
F Details of the Phase diagram
In this section, we expand further on the phase diagrams [Figs. 4(a) and 4(b)] presented in the main text and discuss its context within the broader window of model parameters.
1 Different regimes of interaction strength
We start our discussion from a two-orbital Hubbard model. The interaction effects in these two orbitals can be very different. The low-energy physics in correlated systems depends on the relative strength of the interaction when compared to various energy scales. Briefly speaking, there are four possible cases. (i) When the interaction strength $U$ is infinitesimally small compared to the width of the flat band crossing the Fermi level (the active band), the low-energy physics is mostly described by a weakly interacting Fermi liquid theory. (ii) When $U$ is comparable to the active band width but still rather small compare to the band gap, the interaction effect will dominate, although the inter-band effect is still considered as irrelevant. For example, fractional Chern insulator states are usually considered in this regime. (iii) When the interaction is further increased to be comparable with the overall band widths of both the two top bands, a multi-orbital description becomes necessary. (iv) When the interaction is much stronger than the kinetic energy of the two top-most bands, the two-orbital description itself may fail. This typically requires the value of $U$ to be more than $\sim 2$ times larger than the band widths of these bands considered together. Different interaction strength regimes are summarized in Fig. S2.
Figure S2: Schematic portrayal of the four interaction regimes. Here $D_{12}\sim D_{\rm top}+D_{\rm bottom}+\Delta$, in which $D_{\rm top}\leavevmode\nobreak\ (D_{\rm bottom})$ stands for the band width of the top (bottom) band, and $\Delta$ is the band gap.
2 Filling factor and correlation effects
In multi-orbital interacting systems, the orbital with the smaller band width is more likely to be affected by the on-site interaction. Additionally, the interaction effect will also be affected by the filling factor of each orbital [45, 46]. In Fig. S3, we provide the quasiparticle weight $Z$ as a function of filling factor $\nu=1+x$ in a single-orbital Hubbard model by $U(1)$ slave spin approach [66]. It can be shown that the orbital with filling factor closer to half-filling will experience a stronger correlation effect. In this situation, the effective description of the system is the Anderson lattice model. The orbitals with stronger correlation effects could be treated as the $f$ orbital, and the other can be treated as a conduction band.
Figure S3: Doping level $x=\nu-1$ versus quasiparticle weight $Z$ for one-band Hubbard model solved via $U(1)$ slave spin method. Here we use a 2D square lattice, with interaction strength $U/t=12.5$.
3 Role of hybridization
The hybridization between these two orbitals $V_{\rm hyb}$ plays an important role in the quantum phases in this Hamiltonian. The Kondo coupling between the conduction electrons and the $f$ orbital spin moment is induced by the hybridization through a second order perturbation, which has the form of $J_{K}\sim V_{\rm hyb}^{2}\left(\frac{1}{\epsilon_{f}}+\frac{1}{-\epsilon_{f}+U}\right)$. The energy scale of the Kondo screening strength is depicted by the Kondo temperature $T_{K}\sim\rho_{0}^{-1}\exp(-1/J_{K}\rho_{0})$. In the mean time, the conduction electrons can also mediate the RKKY exchange interaction between the $f$-orbital spin moments in different unit cells, whose strength is controlled by $I\sim\rho_{0}J_{K}^{2}$, in which $\rho_{0}$ is the density of states of the conduction electrons at Fermi level.
Since the Kondo temperature $T_{K}$ and the RKKY exchange strength are controlled by the $V_{\rm hyb}$ in a different manner, changing the strength of $V_{\rm hyb}$ will also change the relative strength between these two quantities, and therefore control the competition between them. As depicted in Fig. 4(a), with a weaker hybridization, the RKKY interaction dominates, and the $f$-orbital spin moments tend to form magnetic ordered state. With a stronger hybridization, the effect of Kondo screening also becomes stronger, and could drive the system across a quantum critical point into a Fermi liquid phase. The spin moment fluctuation near the quantum critical point can lead to superconducting instability at low temperatures.
In the C-EDMFT treatment of the Anderson lattice model, the RKKY interaction is considered as an independent parameter to the hybridization strength, as shown in Eq. (2) in the main text. The quantum phase transition can also be achieved by tuning the value of $I$ with fixed $V_{\rm hyb}$, as demonstrated in Fig. 4(b) in the main text. This is more convenient in the framework of C-EDMFT technologically. The phase with stronger $V_{\rm hyb}$ in Fig. 4(a), which corresponds to the Kondo screened phase, is the same phase with weak exchange interaction $I$ in Fig. 4(b).
We also comment on the effect of on-site interaction strength $U$ in the phase diagram. As mentioned in the previous paragraphs, one of the most important energy scales, the Kondo coupling $J_{K}$, can be influenced by the interaction strength. Since superconductivity arises from spin fluctuations, the transition temperature $T_{c}$ may also be affected. However, in all our calculations, we find $T_{c}/T_{K}$ to be on the order of a few percent across different interaction strengths. In our case, the value of $\epsilon_{f}$ (measured relative to the chemical potential) is relatively small compared to $U$, the effect of $U$ on $J_{K}$ is not too strong, suggesting that $T_{c}$ might not be highly sensitive to $U$.
G Robustness of the superconductivity based on additional numerical results
In this section, we present our additional numerical results. Here, we use the tight-binding model with parameters $(\tilde{v},w,\varepsilon_{D},\theta)=(13{\rm\,meV},15{\rm\,meV},15{\rm\,meV},3.65^{\circ})$. As we have already explained before, the top two moiré bands can be Wannierized with these parameters choice. The band structure and the corresponding orbitals are summarized in Fig. S4.
In addition, using the tight-binding parameters listed in Table S2, we estimate the average hybridization between the two orbitals to be around $V_{\rm hyb}=7\rm\,meV$. We solved this model using C-EDMFT with different interaction strengths $U=40\,\rm meV$ and $U=80\,\rm meV$. Since the interaction strength $U$ is noticeably larger than the value of $\epsilon_{f}-\mu$, we expect that the Kondo coupling is not very sensitive to the value of $U$. The numerical simulation shows that the bare Kondo temperature scales in these two cases are given by $k_{B}T^{0}_{K}=18.43\rm\,meV$ and $k_{B}T^{0}_{K}=18.36\rm\,meV$, respectively. The superconducting transition temperature, and the electronic order transition temperature are also solved, which are shown in Fig. S5. Similar to the results shown in the main text, $T_{c}$ is able to approach a few percent of the Kondo temperature, which indeed aligns with the experimental observation. In addition, since the effective Kondo coupling scale is not very sensitive to the $U$ once it is already in the suitable regime, the transition temperatures are also close to each other, although the interaction strengths are different. This result also demonstrates the robustness of the mechanism we have proposed for superconductivity against interaction parameter tuning.
Figure S4: (a-b) The fat band plots of the two Wannier orbitals. (c) The real-space density distributions of these two orbitals. (d) The density of states or the two-orbital tight-binding model. (e) The density of states without the hybridization between the two orbitals. The band structure parameters are chosen as $(\tilde{v},w,\varepsilon_{D},\theta)=(13{\rm\,meV},15{\rm\,meV},15{\rm\,meV},3.65^{\circ})$.
Figure S5: The superconducting transition temperature $T_{c}$, and the magnetic ordering temperature $T_{\rm order}$ scales as functions of the exchange coupling $I$ with interaction strength $U=40\,\rm meV$ and $U=80\,\rm meV$.
H Connection with experiments
The superconductivity phase in twisted bilayer $\rm WSe_{2}$ has already been observed in two independent experiments [15, 16]. The filling factor can be controlled by the gate voltages, and the superconducting state is discovered when the sample is doped by one hole per moiré unit cell. A correlated phase is also observed in the vicinity of this superconducting phase.
Experiments have provided evidence for the strongly correlated nature of the normal state. The resistivity of Landau Fermi liquid behaves like $\rho(T)\approx\rho_{0}+AT^{2}$ at low temperatures, in which the coefficient $A$ is proportional to the square of effective mass. In tWSe2, when the filling factor approaches the superconducting regime in the phase diagram, the value of $A$ is strongly increased, as shown in Fig. S6: $\sqrt{A}$ increases by more than a factor of $10$. This is highly reminiscent to what is observed in quantum critical heavy fermion metals. By comparison with the latter, it suggests that the normal state of the superconducting tWSe2 is in proximity to a quantum critical point, as we have proposed theoretically.
Figure S6: The coefficient $A$ as a function of filling factor in the experimental observation. The quantity $A$ is defined as the coefficient of $T^{2}$-term in the temperature dependent resistivity in Fermi liquid $\rho(T)\approx\rho_{0}+AT^{2}$. This figure is adapted from Ref. [15].